OpenStudy (anonymous):

Find the difference of (-4x2 - 7x3 - 2) - (8x3 + 3x - 7x2). Be sure to show all of your work in a neat and organized fashion.

OpenStudy (anonymous):

please help i'm confued with this.

OpenStudy (anonymous):

hello?

OpenStudy (callisto):

First, get rid of the bracket by multiplying -1 into the terms inside the bracket. \[(-4x^2 - 7x^3 - 2) - (8x^3 + 3x - 7x^2) = -4x^2 - 7x^3 - 2 - 8x^3 - 3x + 7x^2 \] Then, group the like terms - terms with the same unknown of the same power. \[= -4x^2 - 7x^3 - 2 - 8x^3 - 3x + 7x^2 \]\[=- 7x^3- 8x^3 -4x^2 + 7x^2 - 3x- 2 \]\[=x^3(- 7- 8) +x^2(-4 + 7) - 3x- 2 \] The do the addition, you'll get the answer. Can you work it out?

OpenStudy (anonymous):

thanx, brain freeze